OpenStudy (anonymous):

Fans & Medals!! What is the solution to the system of equations represented by the two equations? Y=3/2x y=-1/2x+4 (__,__)

OpenStudy (anonymous):

You have one value of y already, y = 3/2x Take that an plug it into the second equation, so you get: 3/2x = -1/2x + 4 Solve for x. Can you do that?

OpenStudy (anonymous):

plug in y=3/2x into the bottom equation and solve

OpenStudy (anonymous):

so i would get 3/2=1/2x+4?

OpenStudy (anonymous):

3/2x = -1/2x + 4 Then you need to solve for x.

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

Ok I got 2?

OpenStudy (anonymous):

Yes! So now we know x = 2. Take that and plug it into either of your original equations. You can use either one.

OpenStudy (anonymous):

Although y = 3/2x would be easiest to use.

OpenStudy (anonymous):

yeah, LOLAri, thats the answer of x

OpenStudy (anonymous):

so y= 3 then using 3/2x

OpenStudy (anonymous):

so i will get 2,3

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes